OpenStudy (kainui):

What's the next number in the pattern? 0, 0, 0, ?

OpenStudy (anonymous):

r the number 000

OpenStudy (anonymous):

?

OpenStudy (ikram002p):

interesting ...

OpenStudy (anonymous):

yep

OpenStudy (kainui):

I'll give you a hint, the next number is not 0.

OpenStudy (ikram002p):

im not expert enough thought :P

OpenStudy (anonymous):

\[0, 0, 0, (ノ◕ヮ◕)ノ*:・゚✧ \]

OpenStudy (anonymous):

what r the answer choices

OpenStudy (anonymous):

0 man

OpenStudy (ikram002p):

it could be anything i just need to find anything that doesn't follow ur rule :P

OpenStudy (kainui):

The answer choices are: \[\LARGE answers \in \mathbb{R} \]

OpenStudy (anonymous):

this is weird

OpenStudy (anonymous):

\[0,0,0,\spadesuit, \heartsuit,\diamondsuit, \clubsuit\] is what i get

OpenStudy (camerondoherty):

I can probably bully my math teacher with this c;

OpenStudy (kainui):

Lol in the mean time, check out my own spin on the dirac delta function/kronecker delta!\[\LARGE \delta(x,y)=0^{|x-y|}\] since \[\LARGE 0^0=1\] and for all other powers\[\LARGE 0^p=0\] ok, now back to guessing or should I reveal the answer?

OpenStudy (camerondoherty):

Revalation time c:

OpenStudy (anonymous):

probably 000111222333?

OpenStudy (kainui):

\[\Huge 119\] \[\LARGE f(n)=\frac{(n!)!}{n!}-1\] So n=0, n=1, n=2 all are 0 but n=3 we get 119.

OpenStudy (anonymous):

maybe that is the answer

OpenStudy (camerondoherty):

Duh

OpenStudy (anonymous):

How come I didn't think of that?!

OpenStudy (camerondoherty):

Maybe if you had given us the equation before though :p

OpenStudy (ikram002p):

nice :)

OpenStudy (anonymous):

micw job

ganeshie8 (ganeshie8):

It would be something else for sure had you replied that before ;p

OpenStudy (kainui):

I'll give a medal + fan to someone who can show me a legitimate sequence that has the first 4 or more numbers exactly the same! lol @ganeshie8 I was hoping that no one would have guessed 119 hahaha, I don't know an alternate one this time around haha but you know how tricky I like to be XD

OpenStudy (anonymous):

Here I am supposed to be doing homework and now you're going to make me spend time trying to think of an answer to that question T_T

OpenStudy (ikram002p):

[n/4] floor function bhahahaha

ganeshie8 (ganeshie8):

lets abuse some more familiar functions like logs/exponents/floors/congruences maybe

ganeshie8 (ganeshie8):

Wow! thats really a clever trick.. you can generate a sequence with ANY number of same terms in the start xD

OpenStudy (camerondoherty):

0^5, 0^4, 0^3, 0^2, 0^1, 0^0, 0^-1... Lol, jk

OpenStudy (ikram002p):

yep i also have this one XD [n/4] ( 119 ) xD

OpenStudy (anonymous):

hello

OpenStudy (kainui):

lol nice XD

OpenStudy (ikram002p):

just trolling :P i have infinite of ideas xD

ganeshie8 (ganeshie8):

try finding this : ?, 0,0,0,0,0,0,...

OpenStudy (kainui):

Maybe I need to put restrictions that after the sequence of 4 repeating terms you have only increasing terms after that and no more repeating terms XD but very clever, I'll give you a medal and help you get 99 hahahaha @ikram002p that way you can be not just green, but 99 =P

OpenStudy (ikram002p):

xD so first term is unknown and everything else is 0 or what ?

ganeshie8 (ganeshie8):

yes except for first term, all are 0

OpenStudy (kainui):

does this work for f(1) since after that you multiply by 0?\[\LARGE f(a)=\prod_{n=2}^\infty (a-n)\]

OpenStudy (ikram002p):

[(n)!/(n+1)!] =0 except when n=0 its 1 xD

OpenStudy (camerondoherty):

its 1?

OpenStudy (kainui):

\[\LARGE \frac{(n+1)!}{n!}-n\]

ganeshie8 (ganeshie8):

OMG! that PI function works beautifully XD

ganeshie8 (ganeshie8):

i had \(f(n) = 0^n\) in mind though

ganeshie8 (ganeshie8):

https://oeis.org/A000007

OpenStudy (ikram002p):

wow !

OpenStudy (kainui):

I solved it thinking you just modified my thing slightly XD

OpenStudy (kainui):

I like that 0^n I have heard people say that 0^0 is undefined so this is very interesting to me thanks for showing me @ganeshie8

OpenStudy (ikram002p):

ok how about this one :D 11111110000000123456 ? 00000000000000

ganeshie8 (ganeshie8):

does this work for generating starting 3 zeroes and strictly increasing afterwards \[\rm \large f(n) = \binom{n}{3}\]

OpenStudy (ikram002p):

0^0 could be any number u want :3 i saw 3 pages proof its 17 xD

ganeshie8 (ganeshie8):

we define binomial coefficient for n<k the usual way : \[\dfrac{n(n-1)(n-2)\cdots (n-k+1)}{k!}\]

ganeshie8 (ganeshie8):

we define 0^0 = 1 in number theory (binomial theorem is important for NT) and when working with sequences and series (think of power series of e^x)

ganeshie8 (ganeshie8):

and whenever it makes sense for it to be 1

ganeshie8 (ganeshie8):

thats pretty much what i found so far.. but really its a controversial function

OpenStudy (kainui):

Ooooh fancy I guess it could be considered divergent?

OpenStudy (ikram002p):

yeah its only where u could use it ...

ganeshie8 (ganeshie8):

im not able to find the reply in stackexchange but it argues somethign like this : \[e^x = \sum\limits_{n=0}^\infty \dfrac{x^n}{n!}\]

ganeshie8 (ganeshie8):

if you don't define 0^0 = 1, you will need to explicitly define the value of e^0 the same problem will be there for many other series that require you to evaluate 0^0

OpenStudy (kainui):

Right, that's the only argument I know of too for 0^0=1, it is an interesting function to consider since it's like the identity matrix if your have row m and column n then the entries are just\[\LARGE a_{mn}=0^{|m-n|}\]

ganeshie8 (ganeshie8):

that looks pretty cool !!! when m=n you get 1, and 0 otherwise